OpenStudy (kingbully):

Maria is ordering copy paper for her office. The office is low on white paper and on blue paper. She will pay $3.50 per ream of white paper and $3.75 per ream of blue paper. If the purchase order is for $36.00 and she ordered 10 reams of paper, how many reams of white paper will she receive?

OpenStudy (texaschic101):

w + b = 10 <--- she ordered 10 reams of paper (both white and blue) 3.5w + 3.75b = 36 Those are the equations we will use and we will solve by substitution. w + b = 10 b = -w + 10 so we will sub -w + 10 in for b in the other equation. 3.5w + 3.75b = 36 3.5w + 3.75(-w + 10) = 36 3.5w - 3.75w + 37.50 = 36 3.5w - 3.75w = 36 - 37.50 -.25w = -1.50 w = -1.50/-0.25 w = 6 w + b = 10 6 + b = 10 b = 10 - 6 b = 4 check.. 3.5w + 3.75b = 36 3.5(6) + 3.75(4) = 36 21 + 15 = 36 36 = 36 (correct) she will receive 6 rolls of white paper and 4 rolls of blue paper